OpenStudy (anonymous):

3 Differential equations really don't know how to solve them help plzz 1) dz/dt+ e^(t + z) = 0 2) y' tan x = 7a + y, y(π/3) = 7a, 0 < x < π/2 3) xy'' + 2y' = 16x2 use the substitution u = y'

OpenStudy (dumbcow):

\[\frac{dz}{dt} + e^{t}e^{z} = 0\] \[\frac{dz}{dt} = -e^{t}e^{z} \] \[\frac{dz}{e^{z}} = -e^{t}dt\] integrate both sides \[-\frac{1}{e^{z}} = -e^{t}+C\] \[e^{z} = \frac{1}{e^{t}+C}\] \[z = -\ln (e^{t}+C)\]

OpenStudy (anonymous):

ohh euh thats what i tried but it seems like its not right i dunno why..thanks though

OpenStudy (dumbcow):

\[\frac{dy}{dx}\tan(x) = 7a +y\] \[\frac{dy}{7a+y} = \frac{dx}{\tan(x)}\] integrate both sides \[\ln (7a+y) = \ln (\sin x)+C\] \[7a+y = C \sin(x)\] \[y = C \sin(x) -7a\] \[y(\pi/3) = 7a \rightarrow \frac{\sqrt{3}}{2}C -7a = 7a \rightarrow C = \frac{28\sqrt{3}a}{3}\] \[y = \frac{28\sqrt{3}a}{3} \sin(x) -7a\]

OpenStudy (dumbcow):

#1 is correct http://www.wolframalpha.com/input/?i=dz%2Fdt+%2B+e^t*e^z+%3D+0

OpenStudy (dumbcow):

Letting y' = u, y'' = du/dx \[x \frac{du}{dx} +2u = 16x^{2}\] divide equation by x \[\frac{du}{dx} + \frac{2}{x} u=16x\] This is non-separable so we need to find integrating factor \[\large \rightarrow e^{\int\limits_{}^{}2/x} = e^{2 \ln x} = x^{2}\] multiply equation by x^2 \[ \frac{du}{dx} x^{2} + 2x u=16x^{3}\] \[(x^{2} u)' = 16x^{3}\] integrate both sides \[x^{2} u = 4x^{4} +C\] \[u = 4x^{2}+\frac{C}{x^{2}}\] u = y' = dy/dx \[y = \int\limits_{}^{} 4x^{2}+\frac{C}{x^{2}} dx\] \[y = \frac{4}{3}x^{3} +\frac{C}{x}+K\]